# EEG-Language Modeling for Pathology Detection

### v2

@@ 4 Discussion
The current work presents a first application of multimodal pretraining using natural language and functional brain data in a medical context. Our findings indicate that EEG-language pretraining provides better representations than EEG-only SSL. The richest representations appear to be obtained via a combination of contrastive learning and exposure to a variety of textual information. Specifically, combining information regarding the patient’s clinical history and medication as well as the physician’s description and interpretation of the EEG considerably improved retrieval performance. Such multimodal models were also found to be capable of zero-shot pathology detection. Using linear probing, sizable performance improvements over EEG-only SSL were observed, with the largest gains in contexts with few annotated samples.
Some considerations of the current study deserve mention. Due to clear data privacy concerns, no additional paired EEG-report datasets are publicly available, which currently prevents assessing the generalizability of our results across datasets. Although great care was taken to prevent data leakage and no model development involved any of the evaluation data, future work is required to properly study generalizability and scaling behavior as demonstrated in [22, 11]. While the retrieval analyses suggest that certain models learn richer data representations, a lack of annotations hindered a more detailed assessment of their utility for downstream tasks. Future research could benefit from annotations for specific pathologies, enabling more precise model comparisons. Additionally, we observed lower pathology detection scores for EEG-only SSL than a previous study [8], despite using the same data augmentations. Their work pretrained larger models on multiple datasets to output sequential representations. However, many such adaptations to the EEG encoder or its training could also be applied to EEG-language modeling. Finally, although several models displayed accurate zero-shot pathology detection, the variability in results may be due to the challenges of language modeling with limited data. Further research is needed to explore additional inductive biases or regularization of the text projector to address this issue.
@@ 5 Acknowledgments and Disclosure of Funding
Appendix A Training Details
@@
A.1 Optimization
All models are trained using the LARS optimizer [47] with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only and 0.01 for EEG-language models, scaled with the batch size (BaseLR $\times$ BatchSize/256 [41]). We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v12.3 and PyTorch v1.12.1.
A.2 EEG Encoder
We use a CNN architecture with a residual stream as the EEG encoder for all analyses (Figure 4). The model uses parallel convolutions, involving reflection padding and 1D-convolutions with kernel sizes $\{4,8,16\}$ with 32 filters each. These outputs are concatenated, resulting in a 96 dimensional representation and 747K trainable parameters. We compare input lengths of EEG crops varying from 5 to 60 seconds. This presents a trade-off where longer crops result in a greater information content per crop, while reducing the total sample size. As EEG-only pretraining relies on data augmentations, this introduces an additional influence of crop length. Specifically, longer crop lengths likely make the pretraining task easier, as augmentations introduce relatively lesser distortion due to the greater information content. We therefore compare performance of different crop lengths for both EEG-language and EEG-only pretraining. As the EEG encoder progressively downsamples the signal, we adjust the pooling layers to the input length. These adjustments are shown in Table 4. For EEG-language pretraining we evaluate zero-shot pathology detection, while for EEG-only pretraining we are required to compare the performance of a linear probe. Results are shown in Figure 5. Due to computational resources, we only compare crop lengths for BYOL and ELMl as representations of EEG-only and EEG-language modeling. We observe that for EEG-only pretraining an intermediate crop-length of 20 seconds performs best, which matches the findings by [8]. Meanwhile, zero-shot pathology detection is found to be insensitive to crop lengths of at least 10 seconds, with 60 second crops scoring highest, while the shortest crop length consistently leads to unstable text representations and chance-level performance.
@@ A.5 EEG-Only Pretraining
A.5.1 Data augmentations
For EEG-only pretraining, we adapt the data augmentations proposed by Mohsenvand et al.[8], which were found to perform well on the TUAB dataset. For a given EEG crop, we apply the same augmentation to each channel. Parameters are sampled independently for each EEG crop and uniformly from the ranges displayed in table 6. Augmentations are visualized for a single EEG channel in figure 7.
@@ A.5.1 Data augmentations
Appendix B Details on Data Subsampling
To alleviate class imbalance in the TUEG dataset, we perform data subsampling. We rely on the LLM summaries of reports, which were more consistent in their phrasing regarding the normal or abnormal status. This allowed for a more reliable classification using regular expressions. All reports for which no clear classification was made were omitted. 5015 reports in the potential training set were classified as normal, which were associated with 7526 EEG recordings. For our ‘pretrain’ data subset, we subsampled the abnormal EEGs to match the amount of normal EEG recordings. This resulted in 7526 abnormal EEG recordings, with 6770 reports. Although only a minor subset of these preliminary classifications was manually verified, it is important to note that this process was solely to alleviate severe class imbalance and was not used for further analysis.
For EEG-language modeling, the pretrain subset was effectively smaller, as a report had to be omitted from pretraining when it did not contain at least one relevant heading. Out of the 15144 total EEG files, this resulted in pretrain sample sizes of: 14836 (clinical history cluster), 14320 (medication cluster), 14800 (description cluster), 14794 (interpretation cluster), and 14946 (all clusters).
@@ Appendix C Classification
Appendix D Post-hoc investigation of data leakage
Method Overlap Balanced Accuracy
TS Yes 74.99$\pm$0.86
TS No 74.56$\pm$1.12
ELMe,l All Clusters Yes 82.64$\pm$0.24
ELMe,l All Clusters No 82.28$\pm$0.64
@@ Appendix E Additional visualisations
E.1 EEG embeddings
We provide t-SNE (complexity=40, [52]) visualisations of the averaged EEG embeddings per subject after pretraining. These are post-hoc plots for which we use models trained on the entire pretraining subset and display embeddings of hold-out TUAB patients. Multimodal modeling, specifically for ELMe,l, shows the clearest visual separation between abnormal and normal EEGs, which is in line with the linear probing results.
Figure 9: Example EEG embeddings averaged within-subject of pretrained models on the TUAB hold-out data. The data is projected using t-SNE. The ‘untrained’ and ‘shuffled reports’ plots feature the same setup as the ELMe,l model, with the latter being trained on reports randomly shuffled between subjects.
E.2 Report and content segmentation

### v3

@@ 5 Discussion
The current work presents a first application of multimodal pretraining combining natural language and functional brain data in a medical context. The proposed methodology achieves significantly improved clinical representations compared to EEG-only SSL, while being inexpensive to pretrain. This improvement stems from our novel sub-unit alignment approach in combination with MIL-InfoNCE to address inherent data misalignment challenges. Notably, these multimodal models demonstrated capability for zero-shot pathology detection and label-efficient linear probing with improvements up to 9.7%, an important advancement for clinical applications. We additionally show generalization of ELMs via external validation and clinical event detection tasks.
Some considerations of this study deserve mention. The current limitation of publicly available paired EEG-report datasets presents a challenge for scaling pretraining data. Future work could address this through the generation of synthetic text captions based on clinical metadata. Furthermore, due to computational constraints, analyses on scaling model sizes are yet to be performed. However, we find highly favourable performance of ELMs compared to significantly larger EEG foundation models pretrained on a large number of datasets.
@@ 6 Acknowledgments and Disclosure of Funding
Appendix A Additional Results
A.1 TUEV: Per-Class analysis
We additionally investigated per-class performance as TUEV includes distinctly different event categories (Figure 4). We observe that ELM-MIL scores well across the three clinical events (SPSW, GPED, PLED) with over 3.5% better average scores. However, the model underperformed on artifact and eye movement detection, which may indicate models may lose sensitivity to events not described in the text. Interestingly, a portion of reports include sections on such technical problems, but these were segmented out for the current study. Follow-up research is needed to further investigate the effects of including such text.
Figure 4: Per-class scores for TUEV show that ELM-MIL outperforms for the clinical events. SPSW: Spike and sharp wave, GPED: generalized periodic epileptiform discharges, PLED: periodic lateralized epileptiform discharges.
A.2 Language-independent effects of sub-unit alignment.
Language-independent effects of sub-unit alignment. Given the broad outperforming of ELMs compared to EEG-only models, especially for ELMe,l, we further investigate whether the general setup of multimodal pretraining provides inherent benefits. EEG recordings are split into multiple crops, which in turn are all aligned to the same clinical report during pretraining. It follows that EEG crops of a single recording are indirectly aligned to one another to some extent (Figure 1C). We investigated this hypothesis by shuffling reports between patients prior to pretraining. We find that while embeddings of single EEG crops of an untrained encoder are only minimally more similar within-subject than between-subject (ratio of $\sim$1.1x), this effect is much more pronounced after pretraining ELMe,l on correctly paired reports ($\sim$6.3x), and even more so after pretraining on shuffled reports ($\sim$15.7x; figure 5). Linear probing reveals that training ELMe,l on shuffled reports clearly boosts pathology detection over using an untrained encoder and manages to almost match EEG-only pretraining without the need for augmentations (mean accuracies of 73.70%, 81.04%, 83.69%). On the contrary, the ratios for ELMl are close to 1 after training using paired and shuffled reports, with the latter resulting in decreased pathology detection accuracy.
A.3 Post-hoc investigation of data leakage
Method Overlap Balanced Accuracy
TS Yes 74.99$\pm$0.86
TS No 74.56$\pm$1.12
ELMe,l 4 Clusters Yes 82.64$\pm$0.24
ELMe,l 4 Clusters No 82.28$\pm$0.64
Appendix B Training Details
@@
B.1 Optimization
All models are pretrained using the LARS optimizer [50] with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only, 0.01 for ELMs, and 0.06 for ELM-MIL, scaled with the batch size (BaseLR $\times$ BatchSize/256; [41]). We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v11.3 and PyTorch v1.12.1.
B.2 EEG Encoder
We use a CNN architecture with a residual stream as the EEG encoder for all analyses (Figure 6). The model uses parallel convolutions, involving reflection padding and 1D-convolutions with kernel sizes $\{4,8,16\}$ with 32 filters each. These outputs are concatenated, resulting in a 96 dimensional representation and 747K trainable parameters. We compare input lengths of EEG crops varying from 5 to 60 seconds. This presents a trade-off where longer crops result in a greater information content per crop, while reducing the total sample size. As EEG-only pretraining relies on data augmentations, this introduces an additional influence of crop length. Specifically, longer crop lengths likely make the pretraining task easier, as augmentations introduce relatively lesser distortion due to the greater information content. We therefore compare performance of different crop lengths for both EEG-language and EEG-only pretraining. As the EEG encoder progressively downsamples the signal, we adjust the pooling layers to the input length. These adjustments are shown in Table 8. For EEG-language pretraining we evaluate zero-shot pathology detection, while for EEG-only pretraining we are required to compare the performance of a linear probe. Results are shown in Figure 7. Due to computational resources, we only compare crop lengths for BYOL and ELMl as representations of EEG-only and EEG-language modeling. We observe that for EEG-only pretraining an intermediate crop-length of 20 seconds performs best, which matches the findings by [8]. Meanwhile, zero-shot pathology detection is found to be relatively insensitive to crop lengths of at least 10 seconds, with 60 second crops scoring highest, while the shortest crop length showed unstable learning. For TUSZ and TUEV evaluations, this was solved by lowering the learning rate.
@@ B.5 EEG-Only Pretraining
B.5.1 Data augmentations
For EEG-only pretraining, we adapt the data augmentations proposed by [8], which were found to perform well on the TUAB dataset. For a given EEG crop, we apply the same augmentation to each channel. Parameters are sampled independently for each EEG crop and uniformly from the ranges displayed in table 10. Augmentations are visualized for a single EEG channel in figure 9.
@@ B.5.1 Data augmentations
Appendix C Details on EEG Data
C.1 EEG preprocessing
EEG data received minimal preprocessing (using MNE [54]). First, the initial 10 seconds were removed to reduce the impact of set-up artefacts. Afterwards, a bandpass filter of 0.1-49 Hz was applied and all recordings were resampled to 100 Hz. To reduce the impact of signal artefacts, all EEG signals had their amplitude clipped to $\pm$ 800 $\mu$V. As a large majority of recordings used an average-reference (AR) or linked-ear reference (LE), we only used these recordings and standardized them via transformation to the 20-channel Temporal Central Parasagittal (TCP) montage.
C.2 Subsampling
TUEG contains considerably more abnormal than normal EEGs. As vision-language models have been shown to be sensitive to imbalanced classes [55], we subsample the data to create approximately equal class representation. We rely on the LLM summaries of reports, which were more consistent in their phrasing regarding the normal or abnormal status. This allowed for a more reliable classification using regular expressions. All reports for which no clear classification was made were omitted. 5015 reports in the potential training set were classified as normal, which were associated with 7526 EEG recordings. For our ‘pretrain’ data subset, we subsampled the abnormal EEGs to match the amount of normal EEG recordings. This resulted in 7526 abnormal EEG recordings, with 6770 reports. Although only a minor subset of these preliminary classifications was manually verified, it is important to note that this process was solely to alleviate severe class imbalance and was not used for further analysis.
For EEG-language modeling, the pretrain subset was effectively smaller, as a report had to be omitted from pretraining when it did not contain at least one relevant heading. Out of the 15144 total EEG files, this resulted in pretrain sample sizes of: 14836 (clinical history cluster), 14320 (medication cluster), 14800 (description cluster), 14794 (interpretation cluster), and 14946 (four clusters).
@@ Appendix D Classification
Supervised Learning
For the supervised learning baseline, we use the identical EEG encoder backbone as used for all other analyses and use 60 second crops. We add an MLP (hidden dimensionality of 256) with dropout $p=0.5$ and output dimensionality equal to the amount of classes. The ADAM learning rate is set to 0.001 and we use the validation set to select weight decay out of $[0.1,0.01,0.0001]$. We use a batch size of 256 and train using the cross entropy loss. When using 100% labels, we first train on the training set for up to 50 epochs (with early stopping after 5 epochs without improvement) and select the epoch which resulted in the best validation loss. Subsequently, we continue training on the train and validation sets together until the loss has decreased below the best validation loss.
@@ Appendix E Additional visualisations
E.1 EEG embeddings of pathology
We provide t-SNE (complexity=40, [58]) visualisations of the averaged EEG embeddings per subject after pretraining. These are post-hoc plots for which we use models trained on the entire pretraining subset and display embeddings of hold-out TUAB patients. ELMe,l and ELM-MIL show the clearest visual separation between abnormal and normal EEGs, which is in line with the linear probing results.
Figure 11: Example EEG embeddings averaged within-subject of pretrained models on the TUAB hold-out data (red: abnormal, blue: normal). The data is projected using t-SNE. The ‘untrained’ and ‘shuffled reports’ plots feature the same setup as the ELMe,l model, with the latter being trained on reports randomly shuffled between subjects.
E.2 Within-subject EEG embeddings
We provide additional visualizations of t-SNE projections of EEG crops (Figure 12). Specifically, we compare ELMe,l using InfoNCE and ELM-MIL using MIL-InfoNCE across three temperature parameters $\tau=[0.1,0.3,1.0]$. To do so, we randomly sample three normal (blue shades) and three abnormal (red shades) subjects. We observe that whereas both methods exhibit diminished subject clustering at a higher temperature ($\tau=1.0$), at low temperatures ($\tau=0.1$) this only occurs for InfoNCE. Meanwhile, subject clustering gets more pronounced for MIL-InfoNCE. This may explain the observation that retrieval performance increases by reducing $\tau$ for MIL-InfoNCE, which as a task requires subject rather than class separation per se.
Figure 12: A comparison of subject clustering using t-SNE projections of embeddings of EEG crops. Red (blue) shades indicate three randomly sampled abnormal (normal) subjects.
E.3 Report and content segmentation